please give step by step procedure for finding out solutio of for following matrix using Gauss jordan elimination method with partial pivoting.
3x + 2y - 4z = 3
2x + 3y + 3z = 15
5x - 3y + z = 14

Answer by samuel nkadi(2)



3x+2y-4z=3
2X+3y+3z=`15
5x-3y+z=14
|3 2 -4 |3 |p
|2 3 3 |15 |r
|5 -3 1 |14 | q
we need 1 in the 1st col. of p .we start with row q,with z

1 -3 5 | 14.1 -3 5 |14 | 1 -3 5 | 14
-4 2 3 |3 =0 10 -23|-39|= 0 10 -23 |-39
3 3 2 |15 0 -6 13 |27 | 0 0 4\5|93\5
x=93\5,